OpenStudy (anonymous):

f(x)=3-1/2x in inverse function?

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

so it goes y=3-1/2y 3/2y=3 y=3/(3/2) y=2 f^-1(x)=2?

OpenStudy (anonymous):

@iambatman did I do it right?

OpenStudy (anonymous):

Well no, but is your equation \[f(x) = \frac{ 3-1 }{ 2x }\]

OpenStudy (jhannybean):

I believe it's \[f(x) = 3 - \frac{1}{2}x\]

OpenStudy (anonymous):

Either way it's wrong, your x disappeared somewhere, you let your x = y, instead of switching them.

OpenStudy (jhannybean):

And when @iambatman said replace f(x) with y, he didn't mean replace all x's with y's xD

OpenStudy (jhannybean):

bingo.

OpenStudy (anonymous):

If you could confirm which equation it is, that would be nice :p

OpenStudy (anonymous):

the equation is f(x)=3-(1/2)x

OpenStudy (anonymous):

Oh so what Jhanny said

OpenStudy (anonymous):

Ok lets do this in steps

OpenStudy (anonymous):